OpenStudy (anonymous):

Brad wants to buy flowers for his friend with $45. The daises are $1 each and the roses are $2 each. He buys 3 more daisies than roses. How much did the roses cost?

OpenStudy (anonymous):

@aum I think it meant like how much did he spend on roses total?

OpenStudy (aum):

Let 'x' be the number of roses that he bought. The number of daisies he bought = x + 3 Cost of 'x' roses = x * 2 = 2x cost of (x+3) daisies = (x+3) * 1 = x+3 Total cost = 2x + x + 3 = 3x + 3 = $45 3x = 45 - 3 = 42 x = 42/3 = 14 Amount spent on roses = 14 * 2 = $28

OpenStudy (anonymous):

Thank you! :)